pandasurvive:

Use the sequence to answer the questions. 2.5, 6, 9.5, 13, 16.5, 20, ... What is the recursive rule for the sequence?

pandasurvive:

@Hero when you have a chance.

pandasurvive:

I put A+3.5 but my teacher said it was wrong.

elektroflow:

Another answer is +7 and -3.5

pandasurvive:

how would it be +7?

elektroflow:

+7 first and then subtract 3.5. Try with 2.5 as your first value and you'll get it

pandasurvive:

Would that just be the same as +3.5?

elektroflow:

it is lol but if your teacher says it's wrong, then you should try alternate methods?

pandasurvive:

Lol

elektroflow:

i mean the first thing i got was 3.5 too so... :/

pandasurvive:

lmao maybe I got the right answer but didnt word it right?

elektroflow:

hmm maybe... did you ask your teacher why?

pandasurvive:

she would take 5 days to reply....

elektroflow:

hmm it should be\[X _{a-1}+3.5\]??

pandasurvive:

I didnt put that....

pandasurvive:

i just put A+3.5

elektroflow:

yeah that's very vague..

elektroflow:

you have to state that to find Xn, you have to use the previous term +3.5 to find it

pandasurvive:

Oh ok that makes sense

Hero:

The correct answer to the question depends on the symbology or variables you are expected to use to represent things such as "previous term".

pandasurvive:

Does say what variables.

pandasurvive:

*doesnt @Hero

Hero:

There must be some reference from which the lesson material comes from. Either a textbook or online (or offline) documentation.

pandasurvive:

not that I see any @Hero

pandasurvive:

@elektroflow what did you say it was x a-1 +3.5?

elektroflow:

yeah

pandasurvive:

Thanks

Hero:

Earlier, you used "A" as the variable, so that must be the reference variable for arithmetic series that you were exposed to. It seems that the recursive rule would be \(A_n = A_{n-1} + 3.5\) \(A_n\) represents a particular term of interest. So \(A_1 = 2.5 \\ A_2 = 6 \\ A_3 = 9.5 \\ A_4 = 13 \\ A_5 = 16.5 \\ A_6 = 20\) For example, if we wanted to find \(A_7\) We would write the following: \(A_7 = A_6 + 3.5 = 20 + 3.5 = 23.5\)

Hero:

@pandasurvive

pandasurvive:

Thanks Hero, I think you are the smartest person here tbh.

Hero:

@pandasurvive It's not about who is the smartest. It is about the knower (the student who wants to know), the known (that which the student wishes to know), and the process of knowing (the correct course of action the knower takes in order to master the known).

pandasurvive:

:o God has spoken

ssddss02:

Hero is correct.

ssddss02:

;)